lorena123:

What is the equation in slope-intercept form of the line that passes through the points (−4, 2) and (12, 6)?

lorena123:

help pls

Rylee88:

1) Find the Slope = y2 - y1 / x2 - x1 = 6 - 2 / 12 - -4 = 4/16 = 1/4 This is the slope, m. 2) Now, plug m and one point into: y - y1 = m (x - x1) y - 6 = 1/4 (x - 12) Distribute the 1/4 y - 6 = 1/4 * x - 3 Solve for y y = 1/4*x - 3 + 6 y = 1/4*x + 3 I hope this helps you out, I got this answer from Wyzant.

Rylee88:

I'm not sure if it's right though, I'm clearly not good at these lol.

QuestionCoveBot:

So instead of a confusing way Rylee, did, I'll just simplify it...

Rylee88:

I got it from an app. Not how I do it =)

lorena123:

the answer is y-0.25x+3

lorena123:

y=*

lorena123:

thank u <33

Rylee88:

Thank you who?

Rylee88:

Well I'm taking it, Your welcome ;)